OpenStudy (anonymous):

I think I might've actually answered this right, but I'm not sure. In the triangle below, I'm supposed to find "b", so the side, not the angle. First off I found angle B=119.9 So then, I plugged that into the formula: b=26.4 .sin(119.9)/sin(26.4) I got that b= 54.10 Is this right???

OpenStudy (anonymous):

@ganeshie8 I really tried on this one... Did I use the law of sines correctly?

ganeshie8 (ganeshie8):

finding B is a good start, but after that it looks wrong

ganeshie8 (ganeshie8):

you're using `law of sines` right ?

ganeshie8 (ganeshie8):

Law of sines : \[\large \dfrac{a}{\sin (A)} = \dfrac{b}{\sin (B)} \]

ganeshie8 (ganeshie8):

\[\large \dfrac{24}{\sin (33.7)} = \dfrac{b}{\sin (119.9)} \]

ganeshie8 (ganeshie8):

solve \(\large b\)

OpenStudy (anonymous):

So! I got b=36.56? At first, I was using the data for C where A should be, so I must've done it wrong? Or it could've just been bad all around..

OpenStudy (anonymous):

Also, this is how I put it into wolframalpha http://www.wolframalpha.com/input/?i=24sin%2833.7%29%3Dbsin%28119.9%29

ganeshie8 (ganeshie8):

two mistakes now : 1) you forgot a division sign 2) you need to use degrees

ganeshie8 (ganeshie8):

Here is the correct way put it in wolfram : http://www.wolframalpha.com/input/?i=solve+24%2Fsin%2833.7%29%3Db%2Fsin%28119.9%29&a=TrigRD_D

ganeshie8 (ganeshie8):

b = 37.4979

OpenStudy (anonymous):

Argh. Okay... I almost got it right!! Now I know..

ganeshie8 (ganeshie8):

good :)